### v1

# ORCID: $0$-0003-2317-2297

Abstract
Automatic summarisation has been used efficiently in recent years to condense texts, conversations, audio, code, and various other artefacts. A range of methods, from simple template-based summaries to complex machine learning techniques—and more recently, large language models—have been employed to generate these summaries. Summarising software design patterns is important because it helps developers quickly understand and reuse complex design concepts, thereby improving software maintainability and development efficiency. However, the generation of summaries for software design patterns has not yet been explored. Our approach utilises code features and JavaParser to parse the code and create a JSON representation. Using an NLG library on this JSON representation, we convert it into natural language text that acts as a summary of the code, capturing the contextual information of the design pattern. Our empirical results indicate that the summaries generated by our approach capture the context in which patterns are applied in the codebase. Statistical evaluations demonstrate that our summaries closely align with human-written summaries, as evident from high values in the ROUGE-L, BLEU-4, NIST, and FrugalScore metrics. A follow-up survey further shows that DPS summaries were rated as capturing context better than human-generated summaries.
Keywords:
Design Patterns Code Features Summarisation JSON JavaParser
journal: Empirical Software Engineering
1 Introduction
Software design patterns are reusable solutions to common software design problems kuchana:2004 ; Gamma:1995 . They provide a way for developers to solve problems in a consistent and efficient manner and to communicate their solutions to other developers Nazar:2022 . Design patterns are abstract and conceptual and do not specify how to implement them in a particular programming language or platform Linkedin:2024 . Identifying software design patterns is a well-established research area in computer science and programming, and many studies have been conducted, including Nazar:2022 , Zanoni:2015 , and Thaller:2019 .
Using design patterns in code introduces several challenges. The first challenge is to implement them without a proper understanding of the context or flexibility, which can lead to over-engineering, increased complexity, and reduced code adaptability Profocus:2025 . The second challenge is to ensure a correct and consistent implementation, as factors such as software aging, lack of developer experience, and the mutation of patterns into anti-patterns can pose significant issues Baumgartner:2019 , Kermansaravi:2021 . A third challenge is effectively communicating design patterns to other developers. Although patterns are intended to facilitate communication and collaboration by providing a common vocabulary and understanding of software design, the lack of standardised documentation and varying developer interpretations can result in inconsistent implementations, further complicating effective communication Wedyan:2020 . Therefore, it is crucial to communicate patterns clearly and explicitly using appropriate diagrams, code comments, documentation, and explaining the rationale and benefits of their use.
To address these challenges, we propose the automatic generation of summaries for design patterns, with a focus on the context in which these patterns are used within the source code. The generation of design pattern summaries offers several advantages to software developers. These include enhanced comprehension and communication of design patterns, improved code maintainability, and the facilitation of effective pattern reuse with minimal customisation in their usage contexts. Additionally, generating summaries of design patterns provides concise explanations and clarifies the relationships between patterns Gamma:1995 , thus facilitating communication and decision making between developers working on the same project Linkedin:2024 .
Building on existing study on design pattern detection Nazar:2022 , where they identified software design patterns using source code features, we leverage these features to generate concise descriptions of common design patterns. We refer to this new approach as the Design Pattern Summariser (DPS). To achieve this, we first collected a corpus of Java files from open-source GitHub projects that specifically implement design patterns. In total, we collected ten such projects, which we refer to as the DPS-Corpus. We parsed the DPS-Corpus by constructing an abstract syntax tree (AST) using source code features, ultimately generating a JSON representation. Next, we used an updated pattern detection approach based on the existing work Nazar:2022 to identify design patterns within the DPS-Corpus and subsequently added or updated the respective pattern information in the JSON representation.
Using the SimpleNLG library [^1] https://github.com/simplenlg/simplenlg accessed and verified on 31-03-2025, we converted the JSON information into natural language summary sentences that capture the context in which design patterns are used. To evaluate the efficacy of our approach, we asked human participants to write summaries of the DPS-Corpus files and compared these with the machine-generated summaries produced by our approach. Human summaries were collected through an online survey involving university staff and post-graduate students with expertise in software engineering. We calculated BLEU-4, NIST, ROUGE-L, and FrugalScore statistical metrics to compare the summaries. Additionally, we conducted an online survey to assess the quality of the summaries by asking five quality-related questions. The results indicate that the DPS summaries capture the context of design patterns more effectively than the human summaries.
Our work focuses on summarising nine commonly used software design patterns—comprising three patterns from each category of design patterns: creational, structural, and behavioural. The selected patterns are Abstract Factory, Adapter, Decorator, Façade, Factory Method, Memento, Observer, Singleton, and Visitor.
The main contributions of our work are as follows:
  • We are the first to generate design pattern summaries.
  • We propose a novel code feature-based approach that identifies design patterns and constructs summaries containing contextual information.
  • We have built a corpus of approximately 1,500 Java files for experimentation.
  • An evaluation of our approach against human-generated summaries demonstrates statistically significant results, indicating the strong potential of our method.
  • An anonymous survey reveals that our summaries capture design pattern context more effectively than human-generated summaries.
The paper is organised into the following sections. Section 1 introduces the problem, and Section 2 provides the motivation behind the paper, particularly by discussing what context means. Section 3 discusses the experimental setup, focusing on corpus selection and generation, the code features used, the statistical measures for evaluation, and the research questions. Section 4 describes our three-phase summariser approach—DPS. Section 5 evaluates our work based on statistical measures and survey-based feedback. Section 6 discusses threats to internal and external validity, while Section 7 provides a description of related work in design pattern detection and code summarisation. Section 8 concludes our paper and discusses future directions for this research.
2 Background:Context
McBurney et al. Mcburney:2014 defined context as the purpose of a method within the code in which it exists. This context includes the dependencies of the method and other methods that rely on it. In the context of design patterns, the term refers to the purpose of a class within the pattern—specifically, the role it plays in the class hierarchy, such as whether it acts as a subscriber or a publisher (as in the Observer pattern), as well as the methods invoked within these classes. Although McBurney’s approach focuses solely on the context of the method, we argue that the context of the class is more critical when dealing with design patterns. However, we capture both class and method contexts in our summaries. Thus, we break down the context from the class level to the methods. The following example, taken from the RefactoringGuru GitHub repository, [^2] https://github.com/RefactoringGuru/design-patterns-java/blob/main/src/refactoring_guru/abstract_factory/example/factories/GUIFactory.java accessed and verified on 31-03-2025 illustrates the context.
public interface GUIFactory {
Button create Button ( ) ;
Checkbox create Checkbox ( ) ;
This code example illustrates that a GUI factory produces two types of products, namely Button and Checkbox, as shown in the implementation class below:
public class MacOSFactory implements GUIFactory {
@Override
public Button create Button ( ) {
return new MacOSButton ( ) ;
@Override
public Checkbox create Checkbox ( ) {
return new MacOSCheckbox ( ) ;
The code presented here illustrates how developers can create families of related objects without specifying their concrete classes. The GUIFactory interface establishes a contract for producing two types of products: Button and Checkbox. Concrete factories, such as MacOSFactory, implement this interface to generate platform-specific variants (MacOSButton and MacOSCheckbox). This approach encapsulates the creation logic, enhancing flexibility and scalability. It not only helps developers identify reusable components, but it also supports refactoring efforts, improves maintainability, and promotes the consistent application of design principles across the codebase. Table 1 shows examples of design pattern contextual summaries of the code examples taken from the corpus.
Table 1: Examples of generated summaries by DPS capturing the context
ID Sample Design Pattern Code DPS Summary
1 public class EmailNotificationListener implements EventListener { private String email; public EmailNotificationListener(String email) { this.email = email; } @Override public void update(String eventType, File file) { System.out.println("Email to " + email + ": Someone has performed " + eventType + " operation with the following file: " + file.getName()); } } EmailNotificationListener acts as an observer, which updates String parameter of eventType, File parameter of file and sends notifications. It is a public class that implements EventListener. The only method of EmailNotificationListener is update (void). No methods call for update and update calls no methods.
2 public interface ComputerPart { public void accept(ComputerPartVisitor computerPartVisitor); } ComputerPart acts as a visitor for element ComputerPart, which is inherited by ComputerPartDisplayVisitor, which accepts ComputerPartVisitor parameter of computerPartVisitor. It is a public interface. The only method of ComputerPart is accept (void). No methods call accept and accept calls no methods.
3 public interface GUIFactory { Button createButton(); CheckBox createCheckBox(); } GUIFactory acts as an abstract factory for Button, CheckBox, which is inherited by WindowsOSFactory, MacOSFactory. It is a public interface. The 2 methods of GUIFactory are createButton (Button) and createCheckBox (CheckBox). No methods call createButton and createButton calls no methods. No methods call createCheckBox and createCheckBox calls no methods.
4 public class HtmlDialog extends Dialog { @Override public Button createButton() { return new HtmlButton(); } } HtmlDialog does not have any design pattern. It is a public class that extends Dialog. The only method of HtmlDialog is createButton (Button). No methods call createButton and createButton calls no methods.
5 public class BusAdapter implements Car { private Bus bus; public BusAdapter() { this.bus = new Bus(); } @Override public void drive() { bus.run(); } } BusAdapter acts as an Adapter, which overrides driveCar. It is a public class that implements Car. The only method of BusAdapter is drive (void). No methods call drive and drive calls no methods.
3 Experimental Setup
This section describes the experimental setup of our study, including the construction of the dataset, the code features used for summarisation, the evaluation metrics applied, and the research questions that guide our analysis.
3.1 Data Collection
To construct our corpus, we manually selected ten open-source GitHub repositories using the keyword “design patterns.” These repositories were not general-purpose software projects but were specifically created to illustrate or teach software design patterns. Most of them originate from educational settings (e.g., university course materials) or accompany books and instructional websites – such as the widely used RefactoringGuru repository. While some repositories contained examples for all 23 Gang of Four (GoF) design patterns, others included a broader or narrower set. We manually curated each repository to retain only those directories relevant to the nine design patterns targeted in our study. This process involved selecting example code files from the appropriate directories and discarding test cases, non-Java files, or unrelated examples. The resulting curated dataset, which we refer to as the DPS-Corpus, contains approximately 1,500 Java files.
Table 2 presents the project names, their URLs, and the number of Java files containing pattern details for the nine design patterns that we aim to investigate. The table summarises the distribution of design pattern implementations across the 10 Java projects within our corpus. Each project demonstrates various design patterns, including Abstract Factory, Adapter, Decorator, Facade, Factory Method, Memento, Observer, Singleton, and Visitor. The total number of pattern files per project ranges from 46 to 943, with the project by shihyu having the highest count at 943 files, and the project by shusheng007 having the lowest count at 46 files. Popular patterns such as Singleton and Factory Method are widely implemented, while others like Memento are less common.
Table 2: Projects in the DPS corpus and the number of instances of each design pattern
Project (GitHub path)
Abstract Factory
Adapter
Decorator
Facade
Factory Method
Memento
Observer
Singleton
Visitor
Total
RefactoringGuru
RefactoringGuru/design-patterns-java
9 4 5 8 6 13 5 2 7 59
Iluwatar
iluwatar/java-design-patterns
13 4 4 5 7 3 11 5 8 60
shihyu
shihyu/DesignPatternExample
141 85 114 83 123 57 119 41 180 943
saeidzabardest
saeidzebardast/java-design-patterns
14 4 7 5 5 3 5 1 7 51
JamesZBL
JamesZBL/java_design_patterns
13 4 3 4 7 3 5 4 8 51
Luisburgos
luisburgos/design-patterns
33 11 10 13 19 5 7 2 11 111
shusheng007
shusheng007/design-patterns
6 5 6 5 3 4 6 4 7 46
AbdurRKhalid
AbdurRKhalid/Design-Patterns
10 4 5 8 6 3 5 1 7 49
Quanke
quanke/design-pattern-java-source-code
10 13 10 5 7 3 9 4 8 69
premaseem
premaseem/DesignPatternsJava9
13 7 9 6 10 8 7 5 4 69
3.2 Code Features
Inspired by the work on Design Pattern Detection by Nazar et al. Nazar:2022 , where they employ code features to identify design patterns, we have used code features to build the summaries. In this study, we have increased the number of features from 15 to 20 and renamed some of the features used by Nazar et al. While some features (and their names) remain the same, others, such as NOML, are newly introduced. Additionally, features such as class name (CN) and method name (MN) have been renamed to enhance clarity and understanding. These features can be categorised into basic, object-oriented, and binary features, and further subdivided into class-level and method-level features. Feature extraction is a crucial aspect of code analysis, and accurately identifying a design pattern necessitates thorough code analysis. Table LABEL:tab:features lists the selected feature types, feature abbreviations, names, and descriptions.
The basic features capture the structural and descriptive elements of Java classes, methods, and fields. These include the names, modifiers, data types, and relationships within the source code. The binary features, on the other hand, represent boolean properties of the Java file regardless of its interface or other types, such as isInterfaceOrNot (ION) and isAbstractOrNot (AON), which classify the type of Java file. Object-oriented features such as ExtendsFrom (EXF) and ImplementsFrom (IMF) uncover inheritance and interface implementation details. Object-oriented features further capture inheritance and implementation relationships between classes and interfaces. Such relationships are needed to capture design patterns, as most of the GoF patterns use some sort of inheritance relationship. Together, these features enable a detailed analysis of code structures, dependencies, and design principles, which are the key essence of software design patterns.
These features are particularly valuable for studying and detecting design patterns, as they reflect fundamental object-oriented principles, i.e., encapsulation, inheritance, abstraction, and polymorphism. Analysing features such as method overrides (MethodOverride (MO)), incoming and outgoing method calls (IM, OM), and adherence to class hierarchies (EXF, IMF), helps in identifying the presence of design patterns. Furthermore, metrics such as MethodBodyLineType (MBLT) and NumberOfMethodVariables (NOMV) enable the identification of code smells and refactoring opportunities – useful in identifying any variants of design patterns such as the Observer pattern which is often used in multiple ways such as Subscriber publisher model, as mentioned by Shvets Shvets:2021 , or the commonly used subject observer model. Using NOMV and MBLT, we can determine multiple variants of the same design pattern.
Table 3: The list of features with their types, names (with abbreviations), and descriptions
Type Feature Name Description
Basic ClassName (CN) Name of the class such as WeatherObserver
ClassModifierType (CMT) Access Modifiers for classes (such as public, private etc)
ConstructorModifier (CM) Access Modifier of the constructor (public, private, static, etc)
FieldDataType (FDT) Data type of the field such as int, String etc.
FieldModifierType (FMT) Access Modifier of the field (public, private, static, etc)
MethodName (MN) Name of the method such as updateWeather.
MethodModifierType (MMT) Access Modifier of the method (public, private, static, etc)
MethodReturnType (MRT) The Return Type of a method (void, int, Object etc)
MethodBodyLineType (MBLT) Type of code in a method’s body e.g. assignment statement, condition statements etc.
NumberOfMethodVariables or attributes (NOMV/NOMA) Number of variables/attributes in a method.
NumberOfMethodCalls (NOMC) Number of method calls in a class.
NumberOfMethodLines (NOML) Number of lines in a method.
NumberofIncomingMethods (NIM) Number of calls to this method
NameofIncomingMethods (NaIM) Names of methods that call this method
NumberofOutgoingMethods (NOM) Number of methods a method calls
NameofOutgoingMethods (NaOM) Names of methods a method calls
ConstructorParameters (CP) List of Arguments of the Constructor
MethodParameters (MP) List of Arguments of the Method
MethodOverride (MO) Is Method using @Override or not
IncomingMethods (IM) List of tuples containing Names of Methods called, and their classes
OutgoingMethods (OM) List of tuples containing Names of Methods that call this class, and their classes
Binary isInterfaceOrNot (ION) A binary feature (0/1) if the Java file itself is an interface
isAbstractOrNot (AON) A binary feature (0/1) if the Java file itself is an abstract class
OO EXtendsFrom (EXF) List of classes that the Java file extends from, i.e. inherits
IMplementsFrom (IMF) List of interfaces that the Java file implements
3.3 Evaluation Metrics
We have used four evaluation metrics to verify the efficacy of our approach, namely BLEU-4, NIST, ROUGE-L, and Frugal Score. We describe them one by one in the subsections below.
3.3.1 BLEU-4
The Bilingual Evaluation Understudy (BLEU) metric is widely used in natural language processing. It evaluates the quality of the outcome of generative tasks such as code comment generation Jiang:2017 , pull request descriptions Liu:2019 , and more. BLEU employs N-grams for matching and calculates the ratio of N-gram similarity between generated text and reference texts. In the context of summarisation, it compares an automatically generated summary or translation to one or more reference summaries or translations and calculates the n-gram overlap between them.
The BLEU score is calculated as the geometric mean of modified n-gram precisions, with a brevity penalty to account for overly short candidate summaries. The score is computed as follows:
[…]
Where $P_{n}$ is the ratio of the subsequences with length $n$ in the candidate that are also in the reference. BP is the brevity penalty for short generated sequences, and $\tau_{n}$ is the uniform weight $1/N$. We use corpus level BLEU-4, i.e., $N=4$, as our evaluation metric since it has been demonstrated to be more correlated with human judgments than other evaluation metrics. The higher the BLEU score, the closer the candidate is to the reference. If the candidate is completely identical to the reference, the BLEU score becomes 100%.
For our evaluation, we employed the sentence_bleu function from the nltk package [^3] https://www.nltk.org/_modules/nltk/translate/bleu_score.html accessed and verified on 31-03-2025 to compare the reference and generated summary sentences. The default weights of 0.25 for each n-gram (1-gram to 4-gram) were used to ensure a balanced contribution of different n-gram lengths to the final score. Additionally, we applied the built-in method1 smoothing function to mitigate the issue of overly harsh scores, particularly in cases where non-matching n-grams would otherwise result in a score of 0. Before evaluation, both the reference and generated summary sentences were tokenised to facilitate accurate and consistent scoring.
3.3.2 NIST
NIST (National Institute of Standards and Technology) is a commonly used evaluation metric to assess the quality of automatically generated textual summaries. It compares an automatically generated summary to a reference summary and calculates various scores, such as unigram precision, unigram recall, and unigram F-measure Doddington:2002 . NIST is a variant of BLEU that places more emphasis on rewarding n-grams that are less frequent or more informative. It is calculated as follows:
$\text{NIST}=\sum_{n=1}^{N}\left\{\frac{\sum_{\text{all }n\text{-gram match that reference}}\text{Info}(n\text{-gram})}{\sum_{n\text{-gram reference}}}\right\}\cdot\exp(\beta)\log^{2}\left[\min\left(\frac{|\rho|}{|r|},1\right)\right]$ (2)
Where $\beta$ is chosen to make the brevity penalty factor equal to 0.5 when the number of words in the system output is two thirds of the average number of words in the reference set. The $\rho$ is the average number of words in a reference sentence. Chauhan et al. Chauhan:2023 argue that NIST considers the information gained in each n-gram over the precision of n-grams, which implies that the NIST metric determines the importance of each n-gram in relation to the phrase. The importance of each n-gram is obtained with the help of weights. A correct n-gram will be given high weight if it is less likely to occur or is rare. The length penalty, known as the brevity penalty for the NIST score, is calculated such that small variations in the length of translation sentences do not significantly affect the overall score.
By assigning higher weights to n-grams, NIST encourages systems to generate summaries that contain novel and informative content. Compared to BLEU, the NIST metric does not penalize translation sentences for short length. Like BLEU, NIST is precision-oriented but avoids the pitfall of treating all n-grams equally, leading to a more significant and informative evaluation. The NIST metric (using the nltk.translate.nist_score package [^4] https://www.nltk.org/api/nltk.translate.nist_score.html accessed and verified on 31-03-2024), an enhancement of the BLEU score, was used to assess the informativeness and precision of the summaries. We utilised the sentence_nist function from the nltk.nist package module to compare the reference summaries with the generated summaries. To ensure meaningful and robust comparisons, the n-gram length was set to $n=4$, thus avoiding excessively sparse matches and maintaining a balance between granularity and computational feasibility.
3.3.3 ROUGE-L
The Recall-Oriented Understudy for Gisting Evaluation (ROUGE) metric is commonly used to evaluate the quality of machine-generated text. The ROUGE score compares an automatically generated summary with a reference summary and calculates the overlap between the two. ROUGE primarily measures recall, focusing on how much of the reference summary’s content is captured by the generated summary.
ROUGE-L compares the longest common subsequence (LCS) between the generated summary and the reference summary. Unlike other ROUGE metrics that focus on n-gram overlaps, ROUGE-L captures the sequence of words in the summary that matches the reference, making it particularly useful for capturing the overall structure and fluency of the text. The ROUGE-L score is calculated using the following formula:
[…]
where Precision is the ratio of the length of the LCS to the length of the candidate summary. Recall is the ratio of the length of the LCS to the length of the reference summary. $\beta$ is a parameter that balances the weight between precision and recall. Typically, $\beta$ is set to 1, giving equal weight to precision and recall. Compared to BLEU and NIST, ROUGE-L is particularly strong at capturing the sequence and order of words, which makes it more aligned with human judgment of text coherence and fluency. Additionally, ROUGE-L balances both precision and recall, providing a comprehensive evaluation of how well the generated summary matches the reference in terms of content and structure. By focusing on the LCS, ROUGE-L is better at evaluating summaries that maintain the fluency and logical flow of the original text, which BLEU and NIST might miss.
To calculate the ROUGE-L score, we employed the Python-based ROUGE library [^5] https://pypi.org/project/rouge-score/accessed and verified on 31-03-2025. Within this library, we used the rougeL scorer, which implements the scoring mechanism based on the longest common sequence to evaluate the overlap between generated and reference texts. Additionally, we enabled the built-in stemmer functionality, which removes word suffixes to enhance the accuracy of word matching and improve the robustness of the evaluation. This approach ensures a reliable assessment of text similarity, particularly in terms of fluency and content overlap.
3.3.4 FrugalScore
As the aforementioned metrics are designed for machine translation and measure n-gram similarity, i.e., a summary with similar meaning but differing sequences of n-grams would perform poorly, we employed another measure called FrugalScore as proposed by Phillips et al. Phillips:2022 . FrugalScore is a reference-based metric designed for evaluating Natural Language Generation (NLG) models. It provides a more efficient and cost-effective alternative to traditional metrics while retaining high performance. Compared to BLEU-4, ROUGE-L, and NIST, FrugalScore is more effective for several reasons:
  1. 1.
    Traditional metrics like BLEU-4 and ROUGE-L have been criticised for their low correlation with human judgment. BLEU-4, for instance, was originally designed for machine translation and has been shown to have limitations when applied to other NLP tasks Eddine:2022 . Similarly, ROUGE-L, while useful for summarisation, often fails to capture the significance of human language Blagec:2022 . FrugalScore, on the other hand, has been designed to better align with human evaluators, providing a more accurate reflection of the quality of generated text Eddine:2022 ; Blagec:2022 .
  2. 2.
    BLEU-4 and ROUGE-L were developed with specific tasks in mind: translation and machine summarization, respectively. This makes them less versatile for other NLP tasks. FrugalScore offers a more flexible approach, making it applicable across a wider range of tasks without significant loss of accuracy Blagec:2022 .
  3. 3.
    FrugalScore is designed to be computationally efficient, making it scalable for large datasets and real-time applications. This is a significant advantage over BLEU-4 and ROUGE-L, which can be computationally intensive and less practical for large-scale evaluations Eddine:2022 . Although our corpus is not large compared to other datasets, we still prefer to use FrugalScore to identify if it captures human sentiments in our machine-generated summaries better than other metrics.
  4. 4.
    FrugalScore aims to provide a more holistic evaluation by considering multiple aspects of text quality, such as fluency, coherence, and relevance. This contrasts with BLEU-4 and ROUGE-L, which often focus on n-gram overlap and may miss other important quality dimensions Eddine:2022 ; Blagec:2022 .
The exact formula for FrugalScore can vary based on its implementation and configuration, but a generalized representation is:
[…]
Where EmbeddingSimilarity measures the semantic similarity between sentence embeddings of generated and reference texts. LexicalSimilarity quantifies surface level overlap (e.g., BLEU, ROUGE), and TaskSpecificComponent is an optional term to account for specific evaluation needs (e.g., fluency or diversity). $\alpha$, $\beta$, and $\gamma$ are tunable weights for different components.
To compute FrugalScore, we used the official package available on GitHub [^6] https://github.com/moussaKam/FrugalScore accessed and verified on 31-03-2025. For the evaluation of our results, we employed the mover-score metric. This specific model leverages the BERT-based architecture, a transformer-based language model designed to capture the semantic meaning of text rather than relying solely on n-gram matching. The mover-score component of this metric quantifies semantic similarity using word embeddings to calculate the minimum distance required to align words between the hypothesis and the reference text. Given the computational demands associated with larger language models (LLMs), the medium-sized model was selected to ensure a balance between evaluation accuracy and computational efficiency.
3.4 Research Questions
Our study investigates the following research questions:
  1. 1.
    RQ1: To what extent do automatically generated design pattern summaries align with human-written summaries when evaluated using standard metrics?
    To address this question, we conducted a survey collecting human-written summaries for corpus files and performed a systematic comparison using established evaluation metrics, i.e., BLEU-4, NIST, ROUGE-L, and FrugalScore. The methodology and comparative analysis are detailed in Section 3.3.
  2. 2.
    RQ2: How do human evaluators perceive the quality of DPS summaries compared to human-authored summaries?
    We designed and administered an expert evaluation survey via Prolific [^7] https://www.prolific.com/ accessed and verified on 08/03/2025, an online research platform, to assess subjective quality ratings of both summary types. The experimental setup, participant criteria, and results are presented in Section 5.2
4 Design Pattern Summariser (DPS) - Our Approach
The preliminaries and foundational concepts of our design pattern summariser (DPS) approach have been discussed in Section 3. In the following subsections, we detail the core components/modules of DPS, i.e., parsing the dataset, the design pattern checker, and summary generation modules.
4.1 Parsing
Upon collecting the data and identifying the code features, we construct an Abstract Syntax Tree (AST) based on the corpus. This AST encapsulates key structural elements including class names, method names, and other relevant features. For parsing the code, we employ the JavaParser API [^8] https://javaparser.org/ accessed and verified on 23/04/2025, which processes the input corpus (DPS-Corpus) sequentially, file by file. JavaParser is a robust toolkit for analysing, transforming, and generating Java code. It delivers an AST representation of Java source code, enabling programmatic interaction and manipulation. Compatible with Java versions up to 17, JavaParser is widely adopted for tasks such as code analysis, refactoring, and automated code generation.
The extracted AST data (which contains the feature information as well) is stored in a structured dictionary and consolidated into a JSON file, which serves as the foundation for subsequent summary generation. This JSON representation – depicted as a subset in Figure 1 – constitutes the final output of this module. The figure illustrates an excerpt derived from the AbdurRKhalid [^9] https://github.com/AbdurRKhalid/Design-Patterns accessed and verified on 08/03/2025 ComputerPart file, a component of the Visitor design pattern implementation.
Figure 1: JSON representation of the ComputerPart class from AbdurRKhalid project
4.2 Pattern Checker
This module represents an improved iteration of our prior work, incorporating updated feature sets and heuristics-based algorithms to identify design patterns within the codebase. Unlike the existing version, which relied on a limited set of features, the current (our) approach leverages a broader range of features such as inheritance hierarchies, keyword usage, and other syntactic elements in conjunction with the constructed AST to verify whether the code matches the requisite pattern configuration.
Upon successful identification of a design pattern, the relevant information is embedded into the JSON representation (from the parsing phase). Our methodology aligns with the pattern structures defined by Shvets Shvets:2021 , where we systematically search for specific feature combinations and their corresponding AST representations to validate the presence of the intended pattern.
For example, in the case of the Observer pattern, the algorithm verifies the existence of a subscriber publisher relationship. This entails ensuring that all concrete subscribers inherit from a subscriber interface and override the update method, as required by the pattern’s structural requirements. The process iterates until all essential code structures are confirmed, culminating in the formal identification of the design pattern.
4.3 Summary Generator
We used SimpleNLG [^10] https://github.com/simplenlg/simplenlg, accessed and verified on 23/03/2025, a Java-based natural language generation API, to produce design pattern summaries. This framework leverages lexical resources and realisation components to construct grammatically correct English sentences from structured inputs. The JSON output generated by our Pattern Checker module serves as the input for this phase, with SimpleNLG transforming the structured pattern information into coherent natural language summaries.
These automatically generated summaries effectively capture both the structural implementation and contextual usage of each design pattern within the dataset. The transformation process maintains the semantic relationships present in the original code while presenting them in linguistically appropriate form. Figure 2 illustrates the complete architecture of our approach, demonstrating the integration between pattern identification and summary generation components.
Figure 2: DPS: Our design pattern summarisation approach
5 Empirical Evaluation
As discussed in Section 3.3, to evaluate the efficacy of our approach, we compare our work with the summaries generated by humans.
5.1 RQ1: Statistical Evaluation of DPS against Human Summaries
We adopted a multi-stage approach to gather human written summaries for our study. Initially, we used the Prolific platform, implementing screening questions based on the methodology outlined in Danilova:2021 to collect qualified participants for writing summaries for the dataset. However, this approach proved ineffective, as most respondents either failed to meet our participation criteria or provided non-serious responses.
Consequently, we refined our recruitment strategy by targeting subject matter experts. We distributed our repository to postgraduate students and faculty members specialising in software engineering at Ada University (Azerbaijan). Additionally, we engaged faculty members from the University of Melbourne (Australia), securing meaningful participation from several qualified experts in the field. We set a criterion that the participants should have at least 3 years of experience using design patterns, have intermediate familiarity with the Java programming language, and belong to the software engineering major. Table 1 shows five summaries of sample design patterns taken from the DPS-Corpus.
Our study engaged a cohort of 25 participants, consisting of postgraduate computer science students and software engineering faculty members as mentioned, to generate human-written summaries of the corpus. We required each code sample to be summarised by a minimum of three independent participants to ensure reliability. While doing so, we asked participants to identify the design pattern they think the code contains and then summarise it. During the annotation and summarisation of the code, certain disagreements arose among participants regarding the classification of pattern types. For instance, in some cases, participants exhibited a tendency to identify a pattern as an abstract factory, whereas the correct classification was a factory method. Thus, to verify the level of agreement between these participants we calculated Cohen’s Kappa Carletta:1996 score, which was $>$ 0.7, i.e., there is moderate to high level of agreement between these participants.
Participants were given flexibility in how they completed the task – some summarised entire projects, while others focused on individual files. Each file was summarised by up to three participants. From these, we manually selected one representative summary per file for use in our evaluation. The selected summaries were chosen based on clarity, accuracy, and completeness, with preference given to those that best captured the design intent and usage context. On average, each participants summarised between 25-50 files, including both pattern implementations (code containing design patterns) and non-pattern files (such as client files which were part of the package). We compared DPS summaries with human summaries using four statistical measures discussed in Section 3.3. Table 4 illustrates the values of BLEU-4, NIST, ROUGE-L and FrugalScore calculates as a result of comparison between both summaries. We calculate the score of each pattern in the directory and then take the average of it. The table shows the average score for each project
Table 4: BLEU, NIST, ROUGE and FrugalScore scores for each of the projects when compared with DPS and human summaries
Project BLEU-4 NIST ROUGE-L FrugalScore
RefactoringGuru 0.47 6.4 0.3 0.56
Iluwatar 0.34 5.8 0.4 0.46
shihyu 0.56 8.3 0.5 0.45
saeidzabardest 0.3 6.3 0.39 0.49
JamesZBL 0.46 6.6 0.36 0.48
Luisburgos 0.51 8.5 0.41 0.61
shusheng007 0.39 5.5 0.39 0.45
AbdurRKhalid 0.34 6.8 0.44 0.55
Quanke 0.44 7.7 0.5 0.6
premaseem 0.38 7.2 0.48 0.51
Total 0.419 6.91 0.407 0.516
The results demonstrate that DPS summaries achieve moderate to high quality across all measures – notably, FrugalScore exhibited superior performance, showing the highest correlation with human judgments. This finding is particularly encouraging as it validates our system’s ability to capture the contextual application of design patterns, a key objective of our research.
5.2 RQ2: Quality of Summaries
To further evaluate the quality of the DPS summaries, we asked human experts to rate the DPS generated summaries and human written summaries. We created a project on Prolific and shared a Google form that contained the source code, the summaries and five quality questions. The form contains a subset of selected summaries – at least one DPS summary and one human summary for each pattern – in total ten such instances. We asked the following five survey questions:
  1. 1.
    Which summary is more accurate?
  2. 2.
    Which summary is more concise?
  3. 3.
    Which summary content is more adequate?
  4. 4.
    Which summary better captures the code context?
  5. 5.
    Which summary better captures the design patterns?
The participants were given three options to choose from for each of the questions: 1) machine generated summary, 2) human written summary and 3) both. To minimise bias and maintain anonymity, the DPS generated summaries were labeled as Summary A, while the manually written summaries were labeled as Summary B. We set the following screening criteria for the participants: they must be computer science graduates, must be fluent in English, must have a postgraduate computer science degree or higher and at least 3 years of experience in using software design patterns in industrial settings. Around 15 participants responded to the survey and 10 were selected. The participants belonged to different demographics and fulfilled the screening criteria.
Based on the results, we observe that human-written summaries were rated higher in terms of accuracy, conciseness, and adequacy—reflecting the thoughtful effort and domain knowledge of our expert annotators. At the same time, DPS-generated summaries outperformed in two dimensions: capturing the context of the code and effectively conveying the underlying design pattern. These are precisely the aspects our approach aims to automate, and the results are encouraging. Despite being fully automated, DPS is able to match or even surpass human summaries in the structural and contextual dimensions that matter most for design pattern understanding. Table 5 shows the mean average values for each question for both summaries.
Table 5: Comparison of DPS and Human Summaries
DPS Summary Human Summaries
Accurate 3.5 4.5
Concise 4 5
Adequate 4 5
Context captured 4.5 4
Convey DP 4 3
Based on the results, we can conclude that the DPS summaries capture the design pattern context better than the human-generated summaries – having higher average score.
6 Threats to the Validity
The section discusses the internal and external threats to the validity of the study.
6.1 Threats to internal validity
We have identified three threats to the internal validity of the study.
  1. 1.
    The accuracy and reliability of this study are dependent on the correctness and robustness of the JavaParser. Any inherent defects or limitations in the parser may introduce inaccuracies in the representation or analysis of the source code. During the course of this research, a specific issue was identified wherein the parser occasionally failed to import files belonging to the same module, thereby disrupting interactions between files within the module. This issue, occasionally resulting in incomplete summaries for the affected modules, was resolved by implementing a custom SymbolSolver in JavaParser and reverting to a slightly outdated version of the parser to ensure stability and functionality.
  2. 2.
    Furthermore, this study employs JSON representation for parsing and representing data. Errors in the JSON parsing process or inaccuracies in the data representation could compromise the integrity of the analysis and lead to misinterpretation of the results. As noted earlier, the aforementioned issue with the Java parser could also result in incomplete JSON representations. This challenge was mitigated by adopting the aforementioned adjustments to the JavaParser configuration.
  3. 3.
    Lastly, the study uses Natural Language Generation (NLG) as a tool for generating natural language summaries. While functional, the generated summaries are relatively primitive in their current form. The quality of these summaries could potentially be enhanced by leveraging more advanced tools or techniques, such as SWUM (Software Word Usage Model) or grammar-based generators, which may offer more sophisticated and contextually accurate outputs. Future work could explore the integration of such tools to improve the overall quality and readability of the generated summaries
6.2 Threats to external validity
The current study relies on a specific corpus of Java projects, which may limit the relevancy of the findings to other programming languages. A more generalised and diverse corpus, encompassing a wider range of programming languages, could enhance the applicability of the approach. Future work will aim to expand the corpus to include a broader variety of software projects to validate the approach across different languages.
In addition, the projects included in the corpus are all open-source and may not fully reflect the characteristics of proprietary or industrial codebases. Open-source projects often differ in coding style, documentation practices, and project structure compared to closed-source software. Furthermore, the corpus may underrepresent certain styles of Java development, such as Android applications or large-scale enterprise systems that use frameworks and architectures not present in the selected projects. These differences may limit the generalizability of the findings to broader software development contexts.
7 Related Work
In this section, we elaborate on two main threads of related work, including code summarisation and design pattern identification.
7.1 Code Summarisation
A significant amount of research has been conducted on code summarisation, with several new techniques and approaches being proposed to improve the accuracy and usefulness of the generated summaries. Code summarisation aims to generate brief natural language descriptions for source code. Automatic code summarisation approaches vary from manually crafted templates Sridhara:2010 ; Sridhara:2011 ; Mcburney:2014 , Information retrieval (IR) based Haiduc:2010 ; Haiduc2010a ; Wong:2013 and learning based approaches Allamanis:2016 ; Iyer:2016 ; Movshovitz:2013 .
Creating manually crafted templates to generate code comments is one of the most common code summarisation approaches. Sridhara et al. Sridhara:2010 used a Software Word Usage Model (SWUM) to create a rule-based model that generates natural language descriptions for Java methods. Moreno et al. Moreno:2013 predefined heuristic rules to select information and generate comments for Java classes by combining them. These rule-based approaches have been expanded to cover special types of code artefacts such as test cases Zhang:2011 and code changes Buse:2010 . In general, these templates synthesise comments by extracting keywords from a given source code.
IR approaches are widely used in summary generation and usually search comments from similar code snippets. Haiduc et al. Haiduc2010a apply the Vector Space Model (VSM) and Latent Semantic Indexing (LSI) to generate term-based comments for classes and methods. Their work was replicated and expanded by Eddy et al. Eddy:2013 who exploited a hierarchical topic model. Wong et al. Wong:2015 applied code clone detection techniques to find similar code snippets using code comments. This work is similar to their previous work AutoComment Wong:2013 , which mines human-written descriptions for automatic comment generation from Stack Overflow.
In past years, some studies have attempted to provide natural language summaries using deep learning approaches. Iyer et al. Iyer:2016 presented RNNs with attention to produce summaries that describe C# code snippets and SQL queries. It takes the source code as plain text and models the conditional distribution of the summary. Allamanis et al. Allamanis:2016 applied a neural convolutional attentional model to the problem of summarising the source code snippets into short, name-like summaries. Hu et al. Hu:2018 leverage the API sequence within the source code to assist in the generation of code comments. Their approach TL-CodeSum fuses the source code and API sequences within it for better comment generation. Nazar et al. Nazar:2016 identified summary sentences for code snippets from the web. They treated code snippets on the web as a natural language text and utilised textual features and machine learning ensembles to determine if the specific line in the code snippets could be a summary sentence. They incorporated crowdsourcing practices to gather textual features and annotate datasets. Their work was inspired by the earlier work of Ying et al. Ying:2013 , where they generated summaries of code fragments.
7.2 Design Pattern Identification
Identification of software design patterns is a constantly evolving research area, and new techniques are being developed to improve the accuracy and efficacy of identification methods. Nazar et al. Nazar:2022 has identified design patterns by employing source code analysis by building callgraphs and using source code features. They transformed the code into a verbose representation, applying the word2vec algorithm to that representation and using supervised classifiers to determine whether the Java files contain a specific pattern. Their classifier achieved more than 80% Precision and outperformed state-of-the-art systems substantially in terms of Precision. They also used a web-based code annotation tool called CodeLabeller Nazar:2023 to label the corpus for the study, which is useful for supervised learning-based studies.
Relevant studies using code metrics in the identification of design patterns were conducted by Uchiyama et al. Uchiyama:2011 , who presented a software pattern detection approach using software metrics and machine learning techniques. They identified candidates for the roles that compose design patterns by considering machine learning techniques and utilising basic software metrics. In their book, Lanza and Marinescu Lanza:2007 discussed learning from the information extracted from design pattern instances that usually include variant implementations such as number of accessor methods. Fontana et al. Fontana:2011 introduced the microstructures that are considered the building blocks of design patterns. Thaller et al. Thaller:2019 has built a feature map for pattern instances using neural networks.
Many studies developed tools that use source code or its intermediate representation to identify patterns as well as machine learning models to predict patterns. Lucia et al. De:2011 , Tosi et al. Tosi:2009 and Zhang and Liu Zhang:2013 used static and dynamic analysis (or a combination of both) to develop Eclipse plugins that detect design patterns from source code. Moreno and Marcus Moreno:2012 developed a tool, JStereoType, used for detecting low-level patterns (such as classes and interfaces) to determine the design intent of the source code. Zanoni et al. Zanoni:2015 exploited a combination of graph matching and machine learning techniques to implement a tool called MARPLE-DPD. Niere et al. Niere:2002 designed the FUJABA Tool Suite which provides developers with support for detecting design patterns (including their variants) and design smells. Hautam’aki Hautamaki:2005 , in a thesis article, used a pattern-based solution and tool to teach software developers how to use development solutions in a project.
Some techniques (including tool generation) have applied reverse engineering to identify design patterns from source code and UML artefacts. For instance, Lucia et al. De:2011 built a tool using static analysis and applied reverse engineering through visual parsing of diagrams. Other studies such as Thongrak and Vatanawood Thongrak:2014 , Panich and Vatanawood Panich:2016 and Shi and Olsson Nija:2006 also applied reverse engineering techniques on UML class and sequence diagrams to extract design patterns using ontology. Brown Brown:1996 proposed a method for the automatic detection of design patterns by reverse engineering the SmallTalk code. Several other approaches have exploited machine learning to solve the issue of variants. For example, Chihada et al. Chihada:2015 mapped the design pattern detection problem to a learning problem. Their proposed detector is based on learning from the information extracted from design pattern instances, which usually includes variant implementations. Ferenc et al. Ferenc:2005 applied machine learning algorithms to filter false positives from the results of a graph matching phase, thereby providing better precision in the overall output while considering variants. A recent work by Hussain et al. Hussain:2018 leveraged deep learning algorithms for the organisation and selection of DPs based on text categorisation. To reduce the size of training examples for DP detection, Dong et al. Dong:2008 proposed a clustering algorithm on decision tree learning to detect design patterns.
8 Conclusion & Future Works
Automatic summarisation has seen significant advancements in recent years, enabling the concise representation of texts, dialogues, audio, source code, and other similar software artefacts. Techniques for generation have ranged from basic template-based approaches to sophisticated machine learning methods. Despite these advancements, summarising software design patterns remains an underexplored area. In this initial study, we build on existing work on identifying design patterns using code features to create automated summaries of these patterns. The results show that our method effectively captures the context in which design patterns are applied within the code. Statistical analysis confirms that the generated summaries are closely aligned with human-authored ones, as evidenced by strong performance in metrics such as ROUGE, BLEU, NIST, and FrugalScore. A follow-up survey with expert participants shows that the DPS summaries capture the design pattern context better than the human written summaries.
In the future, we plan to experiment with our approach with large language model (LLM) methods using OpenAI or DeepSeek API to build natural language summaries of software design patterns or capture other aspects in code such as quality concerns. In addition, other summaries generation techniques such as SWUM or grammar-based generators could also be explored.
Appendix A Reproducibility
For the purpose of reproducibility of our results, we have released our complete implementation with the annotated reference set and the result files to the public as an open-source project via our online appendix at https://github.com/najamnazar/DPS-Design-Pattern-Summariser. We have collected all projects in the corpus and have committed them to a GitHub repository which is accessible at https://github.com/SamSike/DPS-Corpus.
Acknowledgements.
We would like to extend our gratitude to the individuals who have dedicated their time and effort to writing manual summaries of the code provided in the corpus and participated in the survey on the Prolific platform.